OpenStudy (anonymous):

tossing a fair die is an experiment that can result in any integer number from 1 to 6 with equal probabilities .let X be the number of dots face of a die .compute E(x) and var(x)

OpenStudy (anonymous):

E(x)=3.5 in your head, since it is the average

OpenStudy (anonymous):

\[\frac{1+2+3+4+5+6}{6}=3.5\]

OpenStudy (anonymous):

What about when two dice is thrown? :) Not so much in head now :D

OpenStudy (anonymous):

ffm can compute the variance, but probably need pencil and paper for this one

OpenStudy (anonymous):

\[Var(X) =E(X^2)-(E(X))^2 \]

OpenStudy (anonymous):

Pencil + Paper or simply Mathematica ? ;-)

OpenStudy (anonymous):

you can do it for two dice in your head too, it is 7, most likely roll

OpenStudy (anonymous):

Wow sat! you are amazing :D

OpenStudy (anonymous):

How do you did it?

OpenStudy (anonymous):

@satellite73 Ping! waiting ...

OpenStudy (anonymous):

ok thank u :) now i can computer variance by myself.